Solve the equations by substitution. If the equations are inconsistent and the solution is an empty set, write "no solution".
x - y = 6
x = y + 2

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Replace in the first equation with the thing is equal to in the second equation, namely . Collect like terms ending with the absurdity that . No solution.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Solve the equations by substitution. If the equations are inconsistent and the solution is an empty set, write "no solution".
x - y = 6
x = y + 2
----
Substitute for "x" and solve for "y":
(y+2) - y = 6
2 = 6
-----
That happend because we assumed there was an (x,y) pair
that saisfied both equations.
There is no such point.
---
The equations are inconsistent.
Cheers,
Stan H.
===================